OpenStudy (anonymous):

Sara bought 7 fish. Every month the number of fish she has doubles. After m months she will have F fish, where F = 7⋅2m . How many fish will Sara have after 2 months if she keeps all of them and the fish stay healthy? A)196 B)28 C)20 D)11

OpenStudy (anonymous):

b=28

OpenStudy (anonymous):

how

OpenStudy (hoblos):

"m" is the number of months, so replace m by 2 and solve for F ( which is the number of fish)

OpenStudy (hoblos):

F = 7(2)(2) = ?

OpenStudy (mathstudent55):

Start with 7 fish. After 1 month, 7 * 2 = 14 fish After 1 more month: 14 * 2 = 28 fish

OpenStudy (anonymous):

\[F=7\times2^{2}\] \[F=7\times4\] \[F=28\]